# Enhancing Outlier Knowledge for Few-Shot Out-of-Distribution Detection with Extensible Local Prompts

### v1

@@ Abstract
Out-of-Distribution (OOD) detection, aiming to distinguish outliers from known categories, has gained prominence in practical scenarios. Recently, the advent of vision-language models (VLM) has heightened interest in enhancing OOD detection for VLM through few-shot tuning. However, existing methods mainly focus on optimizing global prompts, ignoring refined utilization of local information with regard to outliers. Motivated by this, we freeze global prompts and introduce a novel coarse-to-fine tuning paradigm to emphasize regional enhancement with local prompts. Our method comprises two integral components: global prompt guided negative augmentation and local prompt enhanced regional regularization. The former utilizes frozen, coarse global prompts as guiding cues to incorporate negative augmentation, thereby leveraging local outlier knowledge. The latter employs trainable local prompts and a regional regularization to capture local information effectively, aiding in outlier identification. We also propose regional-related metric to empower the enrichment of OOD detection. Moreover, since our approach explores enhancing local prompts only, it can be seamlessly integrated with trained global prompts during inference to boost the performance. Comprehensive experiments demonstrate the effectiveness and potential of our method. Notably, our method reduces average FPR95 by 5.17% against state-of-the-art method in 4-shot tuning on challenging ImageNet-1k dataset, even outperforming 16-shot results of previous methods.
@@ 1 Introduction
Out-of-distribution (OOD) detection [1, 2, 3, 4] aims to distinguish outliers, i.e., samples that do not belong to known in-distribution (ID) classes. It is crucial for industries that require a high level of safety [5], such as face recognition [6] and autonomous driving [7]. Most previous methods in the field of OOD detection handle the problem using single-modal methods and concentrate on post-hoc processing [8, 9] or leveraging outliers [10, 11]. However, they suffer from computational inefficiency or consuming data collection. Recently, as the emergence of vision-language models [12, 13] shows promising results in multi-modal tasks including image caption [14], video understanding [15], and so on, it is promising to exploit textual representations to improve performance of OOD detection given a vision-language model [12] as a prior model. Considering that zero-shot undergoes a domain gap between upstream and target distribution while full-tuning may pose threats to the learned representation, it is meaningful to explore few-shot learning for vision-language model in OOD detection, in which the detector is prohibited from real outliers and only has access to several ID images.
@@
The most challenging scene for OOD detection is that one hard OOD sample is similar to a known class on the whole and only has subtle differences locally, which naturally requires the detector to identify outliers through local outlier regions. However, existing research falls short of refining OOD task via rich local information when subtle OOD samples are exhibited in certain regions, as is shown in Fig. 1. Some methods merely focus on utilizing global features only [16], which ignores local features and inevitably brings about coarse description. Others use the same prompts to match both global and local image features [17, 18], so the gap between them may lead to inaccurate local outlier identification. Consequently, it is straightforward that enhancing regional information to empower the model with local outlier knowledge could be significant to OOD detection.
@@
Figure 1: Comparison of prompt learning for OOD detection task. Prompts with global optimization may fail in challenging OOD samples as they are overall similar to ID samples and only have subtle regional differences. For example, cat and tiger are generally similar (blue boxes) and only differ in forehead (red box). Our approach with local outlier knowledge tackles the issue to some extent.
@@
Motivated by the above observations, we focus on how to explicitly exploit regional-related knowledge, i.e., the learned knowledge is determined by the position of regions, about not only known classes but also unknown samples, which has not been explored before. To tackle the issue, we focus on enhancing local features solely and propose the extensible local prompts for fine local information utilization. Concretely, we decompose global and local prompts for respective features, and propose two integral components: global prompt guided negative augmentation and local prompt enhanced regional regularization. Global prompt is frozen to provide guidance for generating negative augmentation, which could facilitate the exploitation of local outlier knowledge. Fine local prompt is designed for regional regularization to refine regional-related local prompts, which captures the fine local information to better distinguish OOD inputs. Moreover, we also propose corresponding regional-related evaluation metrics to leverage the power of local prompts.
@@
Another advantage of our approach is that it orthogonally explores the benefits of local prompts shown in Fig. 1 and uses hand-crafted global prompts, i.e., a photo of a {class}. Hence, it is extensible to numerous global prompt designing methods [18, 19]. Therefore, we can extend it by seamlessly replacing coarse global prompts with trained ones during inference to improve the performance.
@@
Comprehensive experiments demonstrate the effectiveness of the proposed method. Specifically, our results outperform the state-of-the-art methods in both OOD detection (5.17% reduction on FPR95) and ID accuracy evaluation (1.02%) under 4-shot tuning by a large margin, exhibiting the superiority and the potential of enhancing local prompts for OOD detection and ID classification ability. We summarize our contributions as follows:
@@ 2.1 Out-of-distribution detection with vision-language model
Out-of-distribution detection has been widely studied in the past decades [10, 20, 21, 11]. With the emergence of CLIP [12], which learns outstanding visual and textual representations pre-trained on large-scale image-text pairs and displays excellent ability in various downstream tasks, much attention has been paid to transferring vision-language model to OOD detection task [22, 16, 23, 19, 18, 24]. For instance, MCM [16] adopts the visual and textual representations extracted from CLIP encoders and applies softmax with temperature to better separate ID and OOD data. ZOC [23] generates pseudo unseen class labels with additional modules and defines a novel confidence score accordingly. Clipn [19] enhances the discrimination of OOD samples with extra encoder and data [25].
@@ 2.2 Multi-modal few-shot prompt learning
Since vision-language model like CLIP [12] achieves superior performance in numerous image-text tasks, many attempts have been made to explore few-shot tuning to leverage the power of large pre-trained models. There are mainly two lines of approaches, namely Adapter Learning [26, 27, 28, 29, 30] and Prompt Learning [31, 32]. Both of them froze the pre-trained encoders when fine-tuning a few additional parameters.
Prompt learning promotes the performance of vision-language model by enhancing the prompts in both visual [33, 32] and textual [31, 34] aspects. A common practice is setting textual prompts learnable and forwarding them together with visual inputs. CoOp [31] replaces the hand-crafted prompts with contextual learnable embeddings and optimizes them during few-shot training. In our approach, we freeze global prompts and fine-tune local prompts individually.
@@ 2.3 Global and local information
In Vision Transformer [35], global tokens are specifically designed for image classification task [36, 37]. They represent overall characteristics of images. Local tokens are utilized for dense prediction tasks including segmentation [38, 39], object detection [40, 41], and so on. The utilization of local features has been explored as well, such as pyramid features [42] and window attention [43, 44, 45].
In the field of OOD detection, global features are primarily employed and numerous studies have been conducted concerning local information [17, 18, 46]. GL-MCM [17] takes all tokens into consideration and uses the sum of global and local OOD scores to measure the confidence. LoCoOp [18] keeps global ID textual embeddings away from the interference of ID-irrelevant regions. However, all existing approaches take same prompts for all features. By contrast, our method directly enhances OOD detection with ID-related areas and refines local prompts to leverage local outlier knowledge.
@@ 3.1 Problem definition
Formally, out-of-distribution detection can be viewed as a binary classification that the detector has to identify whether the input image is from ID or OOD space. Labeled training data is composed of $\mathcal{D}^{\mathrm{in}}_{\mathrm{train}}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, where $\bm{x}_{i}$ is sampled i.i.d. from joint data distribution space $\mathcal{P}_{\mathcal{XY}^{\mathrm{in}}}$ and $y_{i}\in\mathcal{Y}^{\mathrm{in}}$. $\mathcal{Y}^{\mathrm{in}}=\{1,\cdots,C\}$ is the label space of ID data and $C$ is the number of classes. We transfer the definition of few-shot learning in classification to OOD detection to fine-tune given vision-language model instead of training from scratch. Under few-shot setting, a small proportion of images in each class (e.g., $1$, $4$, or $16$ images) are extracted for training. $\mathcal{Y}^{\mathrm{out}}$ is the label space of OOD data and there is no overlap between ID and OOD label, i.e., $\mathcal{Y}^{\mathrm{in}}\cap\mathcal{Y}^{\mathrm{out}}=\emptyset$. Typically, the model is unable to obtain any OOD data during training process.
@@
3.2 OOD detetion with CLIP
Feature extraction
CLIP [12] is compromised of image encoder $E_{\mathrm{I}}$ and text encoder $E_{\mathrm{T}}$. Given an image $I$ and a text $T$ describing the corresponding image, they are extracted into feature embeddings, respectively. Specifically, text can be hand-crafted templates [12] or learnable context words [34], i.e., $\bm{t}_{c}=[\bm{v}_{1},\cdots,\bm{v}_{L};\bm{v}_{c}]$, where $\bm{v}_{i},(i=1,\cdots,L)$ is the learnable embedding, and $L$ is the length of context words. $\bm{v}_{c}$ is the embedding of class name. Text feature is obtained from text encoder $\bm{t}=E_{\mathrm{T}}(\bm{t}_{c}):\mathbb{R}^{(L+1)\times d}\rightarrow\mathbb{R}^{d}$.
Image is first split into several patches with an additional class token embedding $\bm{x}\in\mathbb{R}^{(N+1)\times d}$. $N$ and $d$ denote the number of local tokens and hidden dimension, respectively. Both global and local visual features are then extracted from image encoder $[\bm{z}^{\mathrm{g}},\bm{z}^{\mathrm{l}}]=E_{\mathrm{I}}(\bm{x}):\mathbb{R}^{(N+1)\times d}\rightarrow\mathbb{R}^{(N+1)\times d}$.
OOD score calculation
The MCM score [16] is defined as the maximum of similarity after $\mathrm{softmax}$ with temperature:
@@ 3.2 OOD detetion with CLIP
and GL-MCM score [17] considers local information. It applies softmax for local features as well and the maximum is selected as local MCM score. Final OOD score is the sum of MCM score and local MCM score:
@@ 4 Methodology
We propose a coarse-to-fine paradigm to strengthen OOD detection with local outlier knowledge. We first present negative augmentation that leverages local and unknown outlier information. Then, we describe local prompt enhancement along with their training regularization and a regional OOD score for better regional information utilization. Finally, we extend our method to trained global prompts to boost the performance. Detailed structures are shown in Fig. 2.
@@ 4.1 Global Prompt Guided Negative Augmentation
As it is costly and not always effective to rely solely on real outliers, one crucial problem for OOD detection is how to leverage imaginary local outlier knowledge to the detector. We propose to synthesize hard negative images to force the model to learn features with strong relation only to simulate outlier situations while avoiding the requirement of real OOD samples. To this end, we apply a simple and straightforward random crop augmentation to empower the detector with unknown (i.e., OOD) information. Specifically, we randomly crop image input $m$ times and use the hand-crafted template with corresponding class name, i.e., $\bm{t}_{c}=$ “a photo of {class}"""" as text inputs to calculate the image-text similarity with global features from image encoder. We then select $m_{1}$/$m_{2}$ images with the largest/least similarity, respectively. Images with the largest similarity can be seen as positive samples and those with the least similarity serve as hard negative samples. Notably, random crop is helpful to leverage local information in that local prompts can learn regional-related outlier information through space translation by means of random crop. Effectiveness of random crop is verified in Sec. 5.3.
Figure 2: Detailed structure of the proposed regional enhancement framework. Our method consists of global prompt guided negative augmentation and local prompt enhanced regional regularization. We froze global prompts to select regional augmented samples and enhance local prompts to learn regional-related representation that helps improve both ID accuracy and OOD detection.
Global prompts can be viewed as coarse guidance for negative augmentation standing for overall representation and are frozen in our framework. Motivated by previous observations[16, 17], we simply use the basic global prompts during training and select augmented images as they are fairly good for extracting global features. Concretely, the global prompts are used to (1) guide the negative augmentation selection, which has been described above; (2) guide the global OOD score calculation for evaluation, as shown in Sec. 4.3.
It is worth emphasizing that our approach is orthogonal to all existing global prompt optimization strategies, i.e., global prompts are built without tuning in our structure [18], so during inference, it is expected that extending the hand-crafted global prompts with carefully designed ones can promote the performance. Concretely, we replace hand-crafted global prompts with trained ones that have the same shape. Note that our main purpose is to decompose global and local prompts and showcase the effectiveness of local outlier enhancement for OOD detection. Therefore, we do not specifically select the template as it is not the focus of the paper.

### v2

# Local-Prompt: Extensible Local Prompts for Few-Shot Out-of-Distribution Detection

@@ Abstract
Out-of-Distribution (OOD) detection, aiming to distinguish outliers from known categories, has gained prominence in practical scenarios. Recently, the advent of vision-language models (VLM) has heightened interest in enhancing OOD detection for VLM through few-shot tuning. However, existing methods mainly focus on optimizing global prompts, ignoring refined utilization of local information with regard to outliers. Motivated by this, we freeze global prompts and introduce Local-Prompt, a novel coarse-to-fine tuning paradigm to emphasize regional enhancement with local prompts. Our method comprises two integral components: global prompt guided negative augmentation and local prompt enhanced regional regularization. The former utilizes frozen, coarse global prompts as guiding cues to incorporate negative augmentation, thereby leveraging local outlier knowledge. The latter employs trainable local prompts and a regional regularization to capture local information effectively, aiding in outlier identification. We also propose regional-related metric to empower the enrichment of OOD detection. Moreover, since our approach explores enhancing local prompts only, it can be seamlessly integrated with trained global prompts during inference to boost the performance. Comprehensive experiments demonstrate the effectiveness and potential of our method. Notably, our method reduces average FPR95 by 5.17% against state-of-the-art method in 4-shot tuning on challenging ImageNet-1k dataset, even outperforming 16-shot results of previous methods. Code is released at https://github.com/AuroraZengfh/Local-Prompt.
@@ 1 Introduction
Out-of-distribution (OOD) detection (Liu et al., 2020; Cheng et al., 2023b; Jaeger et al., 2022; Tao et al., 2022) aims to distinguish outliers, i.e., samples that do not belong to known in-distribution (ID) classes. It is crucial for industries that require a high level of safety, such as face recognition (Lopez-Lopez et al., 2022) and autonomous driving (Filos et al., 2020). Most previous methods in the field of OOD detection handle the problem using single-modal methods and concentrate on post-hoc processing (Hendrycks & Gimpel, 2016; Wang et al., 2022) or leveraging outliers (Hendrycks et al., 2018; Jiang et al., 2024). However, they suffer from computational inefficiency or consuming data collection. Recently, as the emergence of vision-language models (Radford et al., 2021; Li et al., 2021) shows promising results in multi-modal tasks including image caption (Li et al., 2022a), video understanding (Xu et al., 2021), and so on, it is promising to exploit textual representations to improve performance of OOD detection given a vision-language model (Radford et al., 2021) as a prior model. Considering that zero-shot undergoes a domain gap between upstream and target distribution while full-tuning may pose threats to the learned representation, it is meaningful to explore few-shot learning for vision-language model in OOD detection, in which the detector is prohibited from real outliers and only has access to several ID images.
@@
The most challenging scene for OOD detection is that one hard OOD sample is similar to a known class on the whole and only has subtle differences locally, which naturally requires the detector to identify outliers through local outlier regions. However, existing research falls short of refining OOD task via rich local information when subtle OOD samples are exhibited in certain regions, as is shown in Fig. 1. Some methods merely focus on utilizing global features only (Ming et al., 2022) (blue boxes in Fig. 1), which ignores local features (red boxes in Fig. 1) and inevitably brings about coarse description. Others use the same prompts to match both global and local image features (Miyai et al., 2023b; a), so the gap between them may lead to inaccurate local outlier identification. Consequently, it is straightforward that enhancing regional information to empower the model with local outlier knowledge could be significant to OOD detection.
@@
Figure 1: Comparison of prompt learning for OOD detection task. Prompts with global optimization may fail in challenging OOD samples as they are overall similar to ID samples and only have subtle regional differences. For example, cat and tiger are generally similar (blue boxes) and only differ in forehead (red box). Our approach with local outlier knowledge cares about region difference and tackles the issue to some extent.
@@
Motivated by the above observations, we focus on how to explicitly exploit regional-related knowledge, i.e., the learned knowledge is determined by the position of regions, about not only known classes but also unknown samples, which has not been explored before. To tackle the issue, we propose Local-Prompt, which concentrates on enhancing local features solely and introduces extensible local prompts for fine local information utilization. Concretely, we decompose global and local prompts to exploit global and local features, and term prompts that interact with global/local features as global/local prompts. Then we propose two integral components: global prompt guided negative augmentation and local prompt enhanced regional regularization. Global prompt is frozen to provide guidance for generating negative augmentation, which could facilitate the exploitation of local outlier knowledge. Fine local prompt is designed for regional regularization to refine regional-related local prompts, which captures the fine local information to better distinguish OOD inputs. Moreover, we also propose corresponding regional-related evaluation metrics to leverage the power of local prompts for OOD detection.
@@
Another advantage of our approach is that it orthogonally explores the benefits of local prompts shown in Fig. 1 and uses hand-crafted global prompts, i.e., a photo of a {class}. Hence, it is extensible to all existing global prompt designing methods (Miyai et al., 2023a; Wang et al., 2023). Therefore, we can extend it by seamlessly replacing coarse global prompts with trained ones during inference to improve the performance.
@@
Comprehensive experiments demonstrate the effectiveness of the proposed method. Specifically, our results outperform the state-of-the-art methods in both OOD detection (5.17% reduction on FPR95) and ID accuracy evaluation (1.02%) under 4-shot tuning by a large margin, exhibiting the superiority and the potential of enhancing local prompts for OOD detection and ID classification ability. We additionally carry out various analysis on and all quantitative outcomes and qualitative visualization validate effectiveness. We summarize our contributions as follows:
@@ 2.1 Out-of-distribution detection with vision-language model
Out-of-distribution detection, as a crucial part of open environment (Zeng et al., 2024; Zhu et al., 2023b; Cheng et al., 2023b), has been widely studied in the past decades (Hendrycks et al., 2018; Cheng et al., 2023a; Shu et al., 2023; Zhu et al., 2023a; Liu et al., 2024). With the emergence of CLIP (Radford et al., 2021), which learns outstanding visual and textual representations pre-trained on large-scale image-text pairs and displays excellent ability in various downstream tasks, much attention has been paid to transferring vision-language model to OOD detection task (Fort et al., 2021; Ming et al., 2022; Esmaeilpour et al., 2022; Wang et al., 2023; Miyai et al., 2023a; Park et al., 2023; Bai et al., 2024). For instance, MCM (Ming et al., 2022) adopts the visual and textual representations extracted from CLIP encoders and applies softmax with temperature to better separate ID and OOD data. ZOC (Esmaeilpour et al., 2022) generates pseudo unseen class labels with additional modules and defines a novel confidence score accordingly. Clipn (Wang et al., 2023) enhances the discrimination of OOD samples with extra encoder and data (Sharma et al., 2018).
@@ 2.2 Multi-modal few-shot prompt learning
Since vision-language model like CLIP (Radford et al., 2021) achieves superior performance in numerous image-text tasks, many attempts have been made to explore few-shot tuning to leverage the power of large pre-trained models. There are mainly two lines of approaches, namely Adapter Learning (Zhang et al., 2021; Gao et al., 2023; Guo et al., 2023; Udandarao et al., 2023; Zhu et al., 2023c) and Prompt Learning (Zhou et al., 2022b; Bahng et al., 2022). Both of them froze the pre-trained encoders when fine-tuning a few additional parameters.
Prompt learning promotes the performance of vision-language model by enhancing the prompts in both visual (Jia et al., 2022; Bahng et al., 2022) and textual (Zhou et al., 2022b; a) aspects. A common practice is setting textual prompts learnable and forwarding them together with visual inputs. CoOp (Zhou et al., 2022b) replaces the hand-crafted prompts with contextual learnable embeddings and optimizes them during few-shot training. In our approach, we freeze global prompts and fine-tune local prompts individually.
@@ 2.3 Global and local information
In Vision Transformer (Dosovitskiy et al., 2020; Zeng & Yu, ), global tokens are specifically designed for image classification task (Touvron et al., 2021; He et al., 2022). They represent overall characteristics of images. Local tokens are utilized for dense prediction tasks including segmentation (Cheng et al., 2022; Strudel et al., 2021), object detection (Chen et al., 2022; Meng et al., 2021), and so on. The utilization of local features has been explored as well, such as pyramid features (Wang et al., 2021) and window attention (Liu et al., 2021; Li et al., 2022b; Liu et al., 2022).
In the field of OOD detection, global features are primarily employed and numerous studies have been conducted concerning local information (Miyai et al., 2023b; a; Zhang et al., 2023). GL-MCM (Miyai et al., 2023b) takes all tokens into consideration and uses the sum of global and local OOD scores to measure the confidence. LoCoOp (Miyai et al., 2023a) keeps global ID textual embeddings away from the interference of ID-irrelevant regions. However, all existing approaches take same prompts for all features. By contrast, our method directly enhances OOD detection with ID-related areas and refines local prompts to leverage local outlier knowledge.
@@ 3.1 Problem definition
Few-shot out-of-distribution detection. Formally, out-of-distribution detection can be viewed as a binary classification that the detector has to identify whether the input image is from ID or OOD space. Labeled training data is composed of $\mathcal{D}^{\mathrm{in}}_{\mathrm{train}}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, where $\bm{x}_{i}$ is sampled i.i.d. from joint data distribution space $\mathcal{P}_{\mathcal{XY}^{\mathrm{in}}}$ and $y_{i}\in\mathcal{Y}^{\mathrm{in}}$. $\mathcal{Y}^{\mathrm{in}}=\{1,\cdots,C\}$ is the label space of ID data and $C$ is the number of classes. We transfer the definition of few-shot learning in classification to OOD detection to fine-tune given vision-language model instead of training from scratch. Under few-shot setting, a small proportion of images in each class (e.g., $1$, $4$, or $16$ images) are extracted for training. $\mathcal{Y}^{\mathrm{out}}$ is the label space of OOD data and there is no overlap between ID and OOD label, i.e., $\mathcal{Y}^{\mathrm{in}}\cap\mathcal{Y}^{\mathrm{out}}=\emptyset$. Typically, the model is unable to obtain any OOD data during training process.
@@
Prompt learning with CLIP. CLIP (Radford et al., 2021) is compromised of image encoder $E_{\mathrm{I}}$ and text encoder $E_{\mathrm{T}}$. Given an image $I$ and a text $T$ describing the corresponding image, they are extracted into feature embeddings, respectively. In prompt learning, textual prompts can be hand-crafted templates (Radford et al., 2021), e.g., “a photo of {class}"""" or learnable context words (Zhou et al., 2022a), e.g., $\bm{t}_{c}=[\bm{v}_{1},\cdots,\bm{v}_{L};\bm{v}_{c}]$, where $\bm{v}_{i},(i=1,\cdots,L)$ is the learnable embedding, and $L$ is the length of context words. $\bm{v}_{c}~{}(c\in\{1,\cdots,C\})$ is the embedding of class name. Features of prompts are obtained from text encoder $E_{\mathrm{T}}(\bm{t}_{c}):\mathbb{R}^{(L+1)\times d}\rightarrow\mathbb{R}^{d}$. For simple notation, we use $\bm{t}_{c}$$\bm{t}$$\bm{\hat{t}}$ described below to represent features of prompts as text encoder is not the focus of the paper.
Image is first split into several patches with an additional class token embedding $\bm{x}\in\mathbb{R}^{(N+1)\times d}$. $N$ and $d$ denote the number of local tokens and hidden dimension, respectively. Both global and local visual features are then extracted from image encoder $[\bm{z}^{\mathrm{g}},\bm{z}^{\mathrm{l}}]=E_{\mathrm{I}}(\bm{x}):\mathbb{R}^{(N+1)\times d}\rightarrow\mathbb{R}^{(N+1)\times d}$, where global features $\bm{z}^{\mathrm{g}}$ and local features $\bm{z}^{\mathrm{l}}$ represents overall features and patch-wise regional features, respectively.
In this paper, we denote global prompts and local prompts to hand-crafted/learnable prompts that interact with overall local features and fine local features, respectively.
OOD detection with CLIP. The MCM score (Ming et al., 2022) is defined as the maximum of similarity after $\mathrm{softmax}$ with temperature:
@@
and GL-MCM score (Miyai et al., 2023b) considers local information. It applies softmax for local features as well and the maximum is selected as local MCM score. Final OOD score is the sum of MCM score and local MCM score:
@@ 4 Methodology
We propose Local-Prompt, a coarse-to-fine paradigm to strengthen OOD detection with local outlier knowledge. We first present negative augmentation that leverages local and unknown outlier information. Then, we describe local prompt enhancement along with their training regularization and a regional OOD score for better regional information utilization. Finally, we extend our method to trained global prompts to boost the performance. Detailed structures are shown in Fig. 2.
@@ 4.1 Global Prompt Guided Negative Augmentation
As it is costly and not always effective to rely solely on real outliers, one crucial problem for OOD detection is how to leverage imaginary local outlier knowledge to the detector. We propose to synthesize hard negative images to force the model to learn features with strong relation only to simulate outlier situations while avoiding the requirement of real OOD samples. To this end, we apply a simple and straightforward random crop augmentation to empower the detector with unknown (i.e., OOD) information. Specifically, we randomly crop image input $m$ times and use the hand-crafted template with corresponding class name, i.e., $\bm{t}_{c}=$ “a photo of {class}"$,c=\{1,\cdots,C\}$ as text inputs to calculate the image-text similarity with global features from image encoder. We then select $m_{1}$/$m_{2}$ images with the largest/least similarity, respectively. Images with the largest similarity can be seen as positive samples and those with the least similarity serve as hard negative samples. Notably, random crop is helpful to leverage local information in that local prompts can learn regional-related outlier information through space translation by means of random crop. Effectiveness of random crop is verified in Sec. 5.2.
Figure 2: Detailed structure of the proposed Local-Prompt. Our method consists of global prompt guided negative augmentation and local prompt enhanced regional regularization. We froze global prompts to select regional augmented samples and enhance local prompts to learn regional-related representation that helps improve both ID accuracy and OOD detection.
Global prompts can be viewed as coarse guidance for negative augmentation standing for overall representation and are frozen in our framework. Motivated by previous observations(Ming et al., 2022; Miyai et al., 2023b), we simply use the basic global prompts during training and select augmented images as they are fairly good for extracting global features. Concretely, the global prompts are used to (1) guide the negative augmentation selection, which has been described above; (2) guide the global OOD score calculation for evaluation, as shown in Sec. 4.3.
It is worth emphasizing that our approach is orthogonal to all existing global prompt optimization strategies, i.e., global prompts are built without tuning in our structure (Miyai et al., 2023a), so during inference, it is expected that extending the hand-crafted global prompts with carefully designed ones can promote the performance. Concretely, we replace hand-crafted global prompts with trained ones that have the same shape. Note that our main purpose is to decompose global and local prompts and showcase the effectiveness of local outlier enhancement for OOD detection. Therefore, we do not specifically select the template as it is not the focus of the paper.